2NF3(g) + 3H(g). =. N2(g) + 6HF(g)

If 15.008 L of hydrogen gas is consumed at STP, how many moles of HF would be formed? ( Show full working with the formulas)
1
Expert's answer
2020-04-13T02:35:01-0400

2NF3 (g) + 3H2 (g) = N2 (g) + 6HF (g)

n(H2) = PV / RT = 100 × 15.008 / ( 8.314 × 273.15 ) = 0.66 mol

n(HF) = 2 × n(H2) = 0.66 × 2 = 1.32 mol

Answer: 1.32 mol.